If we start with an original tube of bacteria and tube 1 is a 1:10 dilution of the original, tube 2 is a 1:10 dilution of tube 1, and tube 3 is a 1:10 dilution of tube 2, what is the total dilution of tube 3? Show your calculations.
Answer #1

Total dilution of tube 1 is 1/10

Total dilution of tube 2 = dilution of tube 2 × total dilution of tube 1

= 1/10 × 1/10 = 1/100

Total dilution of tube 3 = dilution of tube 3 × total dilution of tube 2

= 1/10 × 1/100 = 1/1000 or 1:1000